# Interaction imbalanced spin-orbit coupled quantum droplets

### v1

@@ Abstract
We explore the ground states and quench dynamics of one-dimensional quantum droplets with spin-orbit coupling (SOC) and an imbalance in intracomponent interactions. A plethora of miscible ground state stripe and standard (i.e., non-modulated) droplets is found depending on the SOC wavenumber and building upon Gaussian to flat-top background for increasing (decreasing) atom number (interactions). Deformations among the states were accompanied by spin population transfer caused by the Rabi coupling and could be controlled by adjusting the interactions or the SOC parameters. When considering a trap, we identified a transition from a bound to a trapped gas many-body state, characterized by a sign change of the chemical potential, which occurred at lower (larger) atom numbers for tighter traps (stronger interactions). The droplet’s breathing motion was accompanied by minor population transfer, and its frequency increases for a larger intracomponent interaction ratio or reaching a maximum at SOC wavenumbers, where the transition from non-modulated flat-top to stripe droplets occurred. We witness droplet fragmentation for abrupt changes in the Rabi coupling while large amplitude quenches of the SOC wavenumber trigger spin-demixing, resulting in constant amplitude but opposite direction motion of untrapped droplets or in-trap out-of-phase oscillating droplets. Our findings have implications for controlled spin-demixing processes of droplets and the excitation of relevant magnetic bound states.
@@ I Introduction
Attractive interacting many-body systems are historically intriguing and challenging to tackle since they are structurally unstable. However, recent advances in cold atoms allow nowadays to enter this regime which, under specific circumstances, can sustain bound states such as quantum droplets [1, 2], bubbles [3, 4] and extreme nonlinear waves, e.g. the Peregrine soliton [5]. Droplets are highly incompressible states which have already been observed in the laboratory using short-range bosonic mixtures [6, 7, 8, 9, 10, 11], as well as long-range dipolar gases [12, 13, 14, 15]. For short-range settings that we are interested in this work, quantum fluctuations, customarily modelled by the Lee-Huang-Yang (LHY) correction term [16], provide the key mechanism to support droplet formation [17, 2]. A frequently employed framework to describe the properties of these configurations, including stationary states [18, 19], collisions [10, 20], collective [21, 22] and nonlinear [23, 18, 24] excitations, refers to a system of suitably extended Gross-Pitaevskii equations (eGPE) taking into account the dimension [25, 26, 27] and system [28] dependent LHY term. Despite pure bosonic systems, droplets can also exist in Bose-Fermi [29, 30, 31] and spin-orbit coupled (SOC) settings [32, 33, 34].
@@
Yet, droplets have been mainly studied by leveraging the fixed intracomponent interaction to density ratio [17, 35] where the system reduces to a single-component one. The breaking of this condition where the two components are prominently distinguishable has been only very recently considered [36, 37, 38, 39, 40], including also lattice [41] and ring [42] geometries where more complex phases have been identified. On the other hand, the inclusion of SOC into the two-component droplet setting brings additional features such as the build-up of stripe droplet configurations [34, 33] and the emergence of spin transfer is prohibited in the absence of Rabi-coupling [43]. Hence, the SOC parameters (wavenumber and Rabi-coupling) may be exploited to demonstrate the controllability of the ensuing unexplored two-component droplet phases. To address these open questions, we employ a two-component bosonic droplet system in one dimension with unequal short-range intracomponent interactions. We aim to examine the interplay of intracomponent repulsion and SOC parameters on the generation of droplet phases, their nonequilibrium dynamics and accompanying spin-demixing processes.
@@
We find a variety of different ground state droplet structures sharing intercomponent miscibility while exhibiting stripe (non-modulated) shapes for large (small) SOC wavenumber. Independently of the latter, the overall droplet background deforms from Gaussian to flat-top for increasing (decreasing) atom number (interaction). Interestingly, a gradual distortion from the stripe to non-modulated Gaussian and afterwards to flat-top droplets occurs for increasing Rabi-coupling. It is also explicated that the intercomponent population imbalance can be controlled through the aforementioned parametric variations, e.g. it increases for either larger interactions or SOC wavenumber, but it vanishes for increasing Rabi-coupling. An external trap enforces Gaussian stripe droplet distributions of increasing size and peak density for larger atom number and, importantly, a transition from a self-bound to a trapped gas reflected in the sign change of the system’s chemical potential. The respective transition region exists at lower (larger) atom numbers for tighter traps (stronger repulsions).
@@
Following a quench of the trap strength, we excite the droplet’s breathing motion, which is accompanied by relatively small population transfer and whose frequency increases for a larger intracomponent interaction ratio. It also becomes maximum as a function of the SOC wavenumber at the transition region from non-modulated flat-top to stripe droplets. Moreover, a sudden change of the Rabi-coupling leads to suppressed spin-mixing and the generation of fragmented moving droplets without the trap or oscillating excited ones in its presence. Additionally, we showcase that for a quench on the SOC wavenumber, the untrapped droplets of each component move in the same (opposite) directions for small (large) post-quench SOC wavenumbers and feature spin-mixing (de-mixing) on top of breathing (delocalization). A similar phenomenology occurs in the presence of a trap, but droplets oscillate out-of-phase (in-phase) in the trap for larger (smaller) quench amplitudes while being immiscible (miscible).
@@ II Imbalanced SOC droplet setting and equations-of-motion
We consider a quasi-1D two-component bosonic system in the presence of SOC characterized by wavenumber $k_{L}$, and Rabi-coupling $\Omega$ between the involved spin-$\uparrow$ and spin-$\downarrow$ states. The atoms in each spin state interact repulsively, i.e., $g_{\uparrow\uparrow}\neq g_{\downarrow\downarrow}>0$ and the coupling among the spin states is attractive $g_{\uparrow\downarrow}<0$. This condition may give rise to a SOC QD in the presence of quantum fluctuations whose effect is taken to first-order via the LHY correction term. The main characteristic of our setup is that the spin states are assumed to feature different intracomponent repulsions ($g_{\uparrow\uparrow}\neq g_{\downarrow\downarrow}$) which in turn result in intercomponent particle imbalance ($N_{\uparrow}\neq N_{\downarrow}$) due to Rabi-coupling. The SOC QD is described by the appropriate set of coupled eGPEs [44, 45, 33, 34] which in dimensionless units read
@@
In a corresponding experiment the two spin states can be emulated, for instance, with the hyperfine states ||$\ket{\uparrow}\equiv\ket{F=1,m_{F}=-1}$ and ||$\ket{\downarrow}\equiv\ket{F=1,m_{F}=0}$ of 39[6, 7, 10], while considering a total number of atoms $\mathcal{N}\sim[10^{4},10^{6}]$. The quasi-1D geometry is achieved via a tight confinement of frequency $\omega_{\perp}=2\pi\times 600$Hz in the transverse directions, which leads to $a_{\perp}=0.657\,\mu$m. Also, across the elongated direction, a box potential using a digital micromirror device can be used [47]. Moreover, the intervals of 3D scattering lengths $a_{\uparrow\uparrow}=2/N[0.1242a_{0},0.3104a_{0}]$ and $a_{\downarrow\downarrow}=2/N[0.1242a_{0},2.4831a_{0}]$, where $a_{0}$ is the Bohr radius, can be used to reach the effective couplings $g_{\uparrow\uparrow}=[0.2,0.5]$ and $g_{\downarrow\downarrow}=[0.2,4]$ allowing for ratios $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}=[1,20]$. Along the same lines, $a_{\uparrow\downarrow}=2/N[-0.0621a_{0},-0.8070a_{0}]$ in order to have $g_{\uparrow\downarrow}=[-0.1,-1.3]$ for $\delta g=0.1$. The dimensionless Rabi-coupling frequency interval $\Omega=[0,30]$ can be realized by tuning the Raman laser intensity [48, 49] in the range $\tilde{\Omega}=2\pi\times[0-18]$kHz, and the SOC wavenumber interval $k_{L}=[0.1,12]$ corresponds to variations of the laser wavelength [48, 50, 51] from $\lambda_{L}=29.19~{}\mu$m to $243.25$nm [48].
To numerically obtain the ground state of the interaction imbalanced SOC droplets described by the time-independent version of Eqs. (1a) and (1b), we use the imaginary time propagation method exploiting a split-step Crank-Nicolson scheme [52, 53, 54]. On the other hand, to investigate the nonequilibrium dynamics of the SOC droplets, we rely on the real-time propagation method. In dimensionless units, the utilized box size has length $L=307$ and features spatial resolution $dx=0.025$, while the time step of the integrator is fixed to $dt=10^{-5}$. For all the simulations to be presented below, we employ a Gaussian initial ansatz for each component, simultaneously imposing the anti-symmetric condition, i.e., ${\psi}_{\uparrow}(x)=-{\psi}_{\downarrow}(-x)$ on the components which facilitates the numerical identification of the SOC droplet ground state [55].
@@ III.1 SOC droplets in free space
We first focus on SOC parameters $\Omega=0.5$, $k_{L}=0.5$ implying the absence of spatially modulated stripe phases since $k_{L}^{2}<\Omega$. Figure 1(a) presents the component density profiles for different $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$, while the remaining system parameters ($N=100$, $\delta g=0.1$) including the SOC ones are held fixed. As it can be seen, upon increasing $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ there are two important structural modifications in the QD distributions. Namely, a gradual transition from flat-top to Gaussian type miscible droplets takes places. This is understood by the fact that a increase in $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ implies a stronger intercomponent attraction for maintaining $\delta g$ constant. This stronger intercomponent attraction is responsible for the spatial localization of the droplets and hence their deformation to Gaussian type configurations which has also been discussed in two-component droplets without SOC [36]. Additionally, to identify the interaction threshold value beyond which the aforementioned transition occurs, we calculate the total kinetic energy of the droplet as a function $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ as depicted in Fig. 2(a). The sharp increase of the kinetic energy beyond $g_{\uparrow\uparrow}/g_{\downarrow\downarrow}\sim 8$ signifies the transition from the flat-top to the Gaussian droplet. Indeed, flat-top droplets being highly incompressible should minize their kinetic energy [28].
Figure 2: (a) Total kinetic energy, $E_{\textrm{ki}}$, with respect to $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ and SOC characteristics $k_{L}=0.5$ and $\Omega=0.5$. The increasing tendency of $E_{\textrm{ki}}$ with a finite rate for larger $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ indicates the deformation of the droplets to a Gaussian shape. (b) Variation of each component atom number as a function of $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ for different $k_{L}$ (see legend). Population transfer occurs from the strongly interacting spin-down to the weakly spin-up components and it is enhanced for larger $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$. The remaining system parameters are the same with Fig. 1.
@@
Similar ground state QD configurations exist also when the SOC parameters become more important, i.e., for $k_{L}^{2}>\Omega$, where stripe droplets build up in both components. To demonstrate this situation we employ a SOC wavenumber $k_{L}=2$ and Rabi-coupling $\Omega=0.5$, while investigating variations of the interaction ratio $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ [Fig. 3(a)] and the total atom number [Fig. 3(b)]. The other system parameters remain fixed. It becomes evident that either decreasing $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ or increasing $N$ a deformation from a Gaussian-type stripe droplet to a flat-top one occurs. Naturally, flat-top droplets can host a larger number of stripes since they are substantially elongated. As in the $k_{L}=0.5$ case, the weakly interacting component has relatively smaller width and larger peak density. Finally, the transfer of atoms takes place from the strongly interacting spin-down to the weakly interacting spin-up component, see Fig.2(b). This population transfer is almost the same both in magnitude and trend with the $k_{L}=0.5$ scenario.
@@
As a next step, we analyze the impact of the Rabi-coupling on the shape and population imbalance of the two-component droplet. Characteristic density profiles of both droplet components are depicted in Fig. 4(a) with respect to the Rabi coupling. It can be seen that a larger $\Omega$ results in the structural modification from the stripe to standard droplets close to the threshold $k_{L}^{2}=\Omega$. Further increasing $\Omega$ enforces a broadening of the non-modulated droplet distributions towards flat-top profiles and simultaneous suppression of the intercomponent population imbalance. Such density deformations upon Rabi-coupling variations have also been reported in single-component droplets [34]. Hence, an increasing $\Omega$ tends to balance the population among the components, which is attributed to the dominance of the Rabi-coupling energy term as compared to the kinetic, mean-field and LHY energy contributions in the droplet.
Moreover, we find that the intercomponent atom transfer from the majority to the minority component is minimized for high $\Omega$ [56]. This can be directly inferred from the spin populations shown in Fig. 4(b) in terms of $\Omega$ for different interactions and $k_{L}=2$. The spin populations remain nearly constant for low $\Omega$, namely before the transition from stripe to standard configurations occurs. Afterwards, they start to be significantly modified and become almost the same for high enough $\Omega$. This process depends on the intracomponent interactions and in particular the $\Omega$ interval where the spin imbalance is enhanced is larger for stronger interactions which are in accordance to the behaviour of a spin-1/2 repulsive gas [43, 57]. Therefore, the interactions can be used as a knob to control the spin imbalance. It is worth noting that a similar behaviour of the spin populations takes place also for other values of the SOC wavenumber $k_{L}$ and especially the imbalance at relatively small $\Omega$ is more prominent for increasing $k_{L}$ (not shown for brevity).

### v2

@@ Abstract
We explore the ground states and quench dynamics of spin-orbit coupled (SOC) one-dimensional two-component quantum droplets featuring intracomponent interaction imbalance. A plethora of miscible ground state stripe and standard flat-top or Gaussian droplets is found depending on the interplay between the SOC wavenumber and interactions. Deformations among these states are accompanied by controllable spin population transfer. Upon considering a trap we identify a transition from a bound to a trapped gas many-body state, captured through a sign change of the chemical potential, which occurs at lower (larger) atom numbers for tighter traps (stronger interactions). The droplets breathing frequency is found to increase for larger intracomponent interaction ratio or reaches a maximum at SOC wavenumbers where the transition from non-modulated flat-top to stripe droplets exists. Dynamical droplet fragmentation occurs for abrupt changes of the Rabi-coupling, while large amplitude quenches of the SOC wavenumber trigger spin-demixed counterpropagating untrapped droplets or in-trap out-of-phase oscillating ones. Our results offer insights into controlled spin-mixing processes in droplets and the potential excitation of magnetic bound states, opening avenues for further research in this field.
@@ I Introduction
Attractive interacting many-body systems are historically intriguing and challenging to tackle since they are structurally unstable. Recent advances in cold atoms allow nowadays to probe this interaction regime which, under specific circumstances, may host bound states such as quantum droplets  [1, 2], bubbles [3, 4] and extreme nonlinear waves, e.g. the Peregrine soliton [5]. Droplets are highly incompressible many-body states which have already been observed in the cold atom laboratories using short-range bosonic mixtures [6, 7, 8, 9, 10, 11], as well as long-range dipolar gases [12, 13, 14, 15]. For short-range settings, that we are interested in this work, quantum fluctuations customarily modelled by the Lee-Huang-Yang (LHY) correction term [16], provide the key mechanism to support droplet formation [17, 2]. A frequently employed framework to describe these self-bound configurations [17, 18, 19], including their collisions [10, 20], collective [21, 22] and nonlinear [23, 24, 25] excitations, refers to a system of suitably extended Gross-Pitaevskii equations (eGPEs). The latter takes into account the dimension [26, 27, 28] and system [29] dependent LHY term. In addition to pure bosonic systems, droplets can also exist in Bose-Fermi [30, 31, 32] and spin-orbit coupled (SOC) attractively interacting settings [33, 34, 35].
@@
Yet, droplets have been mainly studied by leveraging the fixed intracomponent interaction to density ratio [17, 36] where the system reduces to a single-component one (symmetric droplet setting). The breaking of this condition where the two components are prominently distinguishable has been only very recently considered [37, 38, 39, 40, 41], including also lattice [42] and ring [43] geometries where more complex phases have been identified. On the other hand, the inclusion of the SOC contribution into the two-component droplet setting can bring into play additional features such as the build-up of stripe droplet configurations [35, 34] and the emergence of spin transfer, both, being prohibited in the absence of Rabi-coupling [44]. Hence, the SOC parameters (wavenumber and Rabi-coupling) may be exploited to create richer two-component droplet phases and tuned to demonstrate controllability of the ensuing unexplored two-component droplet phases. In this context, for instance, structural deformations from stripe to non-modulated either flat-top or Gaussian-type droplet distributions and vice versa can occur by means of varying the SOC parameters. These processes, that are absent without the SOC term are anticipated to facilitate adjustable intercomponent population transfer. Similarly, they can lead to transitions i) from droplet to gas states and ii) from particle imbalanced droplets to symmetric ones. Another interesting prospect, concerns the dynamical response of SOC droplets. For instance, quenching the appropriate SOC parameters, can give rise to counterpropagating or oscillating droplets or lead to their spontaneous fragmentation.
To address these open questions, we employ a one-dimensional two-component bosonic droplet system featuring a SOC contribution and unequal short-range intracomponent interactions. We aim to examine the highly unexplored interplay of intracomponent repulsion and SOC parameters on the generation of droplet phases, their nonequilibrium dynamics and accompanying spin-demixing processes. Our study covers systems both in free space and under the influence of an external trap.
@@
We find a variety of different ground state droplet structures sharing intercomponent miscibility while exhibiting stripe (non-modulated) shapes for large (small) SOC wavenumber. Independently of the latter, the overall droplet background deforms from Gaussian to flat-top for increasing (decreasing) atom number (interaction) [35]. Interestingly, a gradual distortion from stripe to non-modulated Gaussian droplets and consecutively to flat-top ones occurs for increasing Rabi-coupling an outcome consisting one of the central results of this work. It is also explicated, for the first time, that intercomponent population imbalance can be controlled through the parametric variations of the SOC terms. Particularly, population imbalance is found to increase for larger interactions or SOC wavenumber, but it vanishes for increasing Rabi-coupling. We also showcase that an external trap enforces Gaussian stripe droplet formation of increasing size and peak density for larger atom number and, importantly, a transition from a self-bound to a trapped gas [40] reflected in the sign change of the system’s chemical potential. The respective transition region exists at lower (larger) atom numbers for tighter traps (stronger repulsions).
@@
Following a quench of the trap strength, we excite the droplet’s breathing motion, which is accompanied by relatively small population transfer. We demonstrate that the frequency of this mode increases for a larger intracomponent interaction ratio and becomes maximum for SOC wavenumbers at the transition from non-modulated flat-top to stripe droplets. Such a behavior is in line with earlier findings but referring to symmetric droplets systems [35]. Moreover, sudden changes of the Rabi-coupling are shown to lead to suppressed spin-mixing and the generation of fragmented moving droplets without the trap or oscillating excited ones in its presence. Additionally, we find that for SOC wavenumber quenches, untrapped droplets of individual components propagate in the same (opposite) direction for small (large) post-quench values. They also experience spin-mixing on top of an overall breathing and de-mixing while delocalizing respectively. Harmonically confined droplets feature out-of-phase (in-phase) oscillations for larger (smaller) quench amplitudes while being immiscible (miscible).
@@ II Imbalanced SOC droplet setting and equations-of-motion
We consider a quasi-1D two-component bosonic system in the presence of SOC characterized by wavenumber $k_{L}$, and Rabi-coupling $\Omega$ between the involved spin-$\uparrow$ and spin-$\downarrow$ states. The atoms in each spin state interact repulsively, i.e., $g_{\uparrow\uparrow}\neq g_{\downarrow\downarrow}>0$ and the coupling among the spin states is attractive $g_{\uparrow\downarrow}<0$. This condition may give rise to a SOC QD in the presence of quantum fluctuations whose effect is taken to first-order via the LHY correction term. The main characteristic of our setup is that the spin states are assumed to feature different intracomponent repulsions ($g_{\uparrow\uparrow}\neq g_{\downarrow\downarrow}$) which in turn result in intercomponent particle imbalance ($N_{\uparrow}\neq N_{\downarrow}$) due to Rabi-coupling. The SOC QD is described by the following set of coupled eGPEs [45, 46, 34, 35, 47] which in dimensionless units read
@@
In a corresponding experiment the two spin states can be emulated, for instance, with the hyperfine states ||$\ket{\uparrow}\equiv\ket{F=1,m_{F}=-1}$ and ||$\ket{\downarrow}\equiv\ket{F=1,m_{F}=0}$ of 39[6, 7, 10], while considering a total number of atoms $\mathcal{N}\sim[10^{4},10^{6}]$. The quasi-1D geometry is achieved via a tight confinement of frequency $\omega_{\perp}=2\pi\times 600$Hz in the transverse directions, which leads to $a_{\perp}=0.657\,\mu$m. Also, across the elongated direction, a box potential using a digital micromirror device can be used [48]. Moreover, the intervals of 3D scattering lengths $a_{\uparrow\uparrow}=2/N[0.1242a_{0},0.3104a_{0}]$ and $a_{\downarrow\downarrow}=2/N[0.1242a_{0},2.4831a_{0}]$, where $a_{0}$ is the Bohr radius, can be used to reach the effective couplings $g_{\uparrow\uparrow}=[0.2,0.5]$ and $g_{\downarrow\downarrow}=[0.2,4]$ allowing for ratios $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}=[1,20]$. Along the same lines, $a_{\uparrow\downarrow}=2/N[-0.0621a_{0},-0.8070a_{0}]$ in order to have $g_{\uparrow\downarrow}=[-0.1,-1.3]$ for $\delta g=0.1$. The dimensionless Rabi-coupling frequency interval $\Omega=[0,30]$ can be realized by tuning the Raman laser intensity [49, 50] in the range $\tilde{\Omega}=2\pi\times[0-18]$kHz, and the SOC wavenumber interval $k_{L}=[0.1,12]$ corresponds to variations of the laser wavelength [49, 51, 52] from $\lambda_{L}=29.19~{}\mu$m to $243.25$nm [49]. At this point, it is important to note that the aforementioned variation in $\lambda_{L}$ is related to the range of the non-dimensional $k_{L}$ values that we are considering for a specific transverse trap frequency. The latter can be flexibly adjusted in corresponding quasi-1D experiments [36]. This in turn implies that the actual range of the dimensionful SOC wavenumbers can be accordingly tuned by means of changing the transverse trap strength (and hence $a_{\perp}$) and the angle between the Raman lasers to match the non-dimensional $k_{L}$ parameter used in our simulations.
To numerically obtain the ground state of the interaction imbalanced SOC droplets described by the time-independent version of Eqs. (1a) and (1b), we use the imaginary time propagation method exploiting a split-step Crank-Nicolson scheme [53, 54, 55]. On the other hand, to investigate the nonequilibrium dynamics of the SOC droplets, we rely on the real-time propagation method. In dimensionless units, the utilized box size has length $L=307$ and features spatial resolution $dx=0.025$, while the time step of the integrator is fixed to $dt=10^{-5}$. We remark that by increasing the spatial resolution, i.e. the number of grid points, in our simulations, the observables to be presented below remain un-altered. This ensures the numerical convergence of our results both for the ground state and the dynamics. For all the simulations to be presented below, we employ a Gaussian initial ansatz for each component, simultaneously imposing the anti-symmetric condition, i.e., ${\psi}_{\uparrow}(x)=-{\psi}_{\downarrow}(-x)$ on the components which facilitates the numerical identification of the SOC droplet ground state [56].
@@ III.1 SOC droplets in free space
We first focus on SOC parameters $\Omega=0.5$, $k_{L}=0.5$ implying the absence of spatially modulated stripe phases since $k_{L}^{2}<\Omega$, a condition that was explicated e.g. in Refs. [46, 35, 45, 49]. Figure 1(a) presents the component density profiles for different $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$, while the remaining system parameters ($N=100$, $\delta g=0.1$) including the SOC ones are held fixed. As it can be seen, upon increasing $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ there are two important structural modifications in the QD distributions. Namely, a gradual transition from flat-top to Gaussian type miscible droplets takes places. This is understood by the fact that a increase in $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ implies a stronger intercomponent attraction for maintaining $\delta g$ constant. This stronger intercomponent attraction is responsible for the spatial localization of the droplets and hence their deformation to Gaussian type configurations which has also been discussed in two-component droplets without SOC [37]. Additionally, to identify the interaction threshold value beyond which the aforementioned transition occurs, we calculate the total kinetic energy of the droplet as a function $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ as depicted in Fig. 2(a). The sharp increase of the kinetic energy beyond $g_{\uparrow\uparrow}/g_{\downarrow\downarrow}\sim 8$ signifies the transition from the flat-top to the Gaussian droplet. Indeed, flat-top droplets being highly incompressible should minimize their kinetic energy [29].
Figure 2: (a) Total kinetic energy, $E_{\textrm{ki}}$, with respect to $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ and SOC characteristics $k_{L}=0.5$ and $\Omega=0.5$. The increasing tendency of $E_{\textrm{ki}}$ with a finite rate for larger $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ indicates the deformation of the droplets to a Gaussian shape. The dotted lines provide a guide to the eye for quantifying the intersection of the interaction regions where $E_{\textrm{ki}}$ becomes almost zero (around $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}\approx 1$) and the one where $E_{\textrm{ki}}$ starts to increase (at $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}>>1$). (b) Variation of each component atom number as a function of $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ for different $k_{L}$ (see legend). Population transfer occurs from the strongly interacting spin-down to the weakly spin-up components and it is enhanced for larger $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$. The remaining system parameters are the same with Fig. 1.
@@
Similar ground state QD configurations exist also when the SOC parameters become more important, i.e., for $k_{L}^{2}>\Omega$, where stripe droplets build up in both components. To demonstrate this situation we employ a SOC wavenumber $k_{L}=2$ and Rabi-coupling $\Omega=0.5$, while investigating variations of the interaction ratio $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ [Fig. 3(a)] and the total atom number [Fig. 3(b)]. The other system parameters remain fixed. It becomes evident that either decreasing $g_{\downarrow\downarrow}/g_{\uparrow\uparrow}$ or increasing $N$ a deformation from a Gaussian-type stripe droplet to a flat-top one occurs. Naturally, flat-top droplets can host a larger number of stripes since they are substantially elongated, as was also reported in the respective interaction balance setting [35]. As in the $k_{L}=0.5$ case, the weakly interacting component has relatively smaller width and larger peak density. Finally, the transfer of atoms takes place from the strongly interacting spin-down to the weakly interacting spin-up component, see Fig.2(b). This population transfer is almost the same both in magnitude and trend with the $k_{L}=0.5$ scenario.
@@
As a next step, we analyze the impact of the Rabi-coupling on the shape and population imbalance of the two-component droplet. Characteristic density profiles of both droplet components are depicted in Fig. 4(a) with respect to the Rabi coupling. It can be seen that a larger $\Omega$ results in the structural modification from the stripe to standard droplets close to the threshold $k_{L}^{2}=\Omega$. Further increasing $\Omega$ enforces a broadening of the non-modulated droplet distributions towards flat-top profiles and simultaneous suppression of the intercomponent population imbalance. Such density deformations upon Rabi-coupling variations have also been reported in single-component droplets [35]. Hence, an increasing $\Omega$ tends to balance the population among the components, which is attributed to the dominance of the Rabi-coupling energy term as compared to the kinetic, mean-field and LHY energy contributions in the droplet. It is interesting to notice that other phases, such as the plane-wave and the zero-momentum ones, which have been reported in the context of SOC gases [57] have not been identified in the present setup. Such states could be unveiled in the respective excitation spectrum by considering, for instance, non-zero quasi-momenta. Such a study worths to be pursued in future works.
Moreover, we find that the intercomponent atom transfer from the majority to the minority component is minimized for high $\Omega$ [58]. This can be directly inferred from the spin populations shown in Fig. 4(b) in terms of $\Omega$ for different interactions and $k_{L}=2$. The spin populations remain nearly constant for low $\Omega$, namely before the transition from stripe to standard configurations occurs. Afterwards, they start to be significantly modified and become almost the same for high enough $\Omega$. This process depends on the intracomponent interactions and in particular the $\Omega$ interval where the spin imbalance is enhanced is larger for stronger interactions. This feature is in accordance to the behaviour of a spin-1/2 repulsive gas [44, 59] where an enhanced Rabi-coupling favors a superposition of spin-balanced components. Therefore, the interactions can be used as a knob to control the spin imbalance. Recall also that such spin transfer processes and in general imbalanced mixed droplet phases are absent for corresponding intracomponent interaction symmetric settings as was shown e.g. in Refs. [46, 34, 35]. It is worth noting that a similar behaviour of the spin populations takes place also for other values of the SOC wavenumber $k_{L}$ and especially the imbalance at relatively small $\Omega$ is more prominent for increasing $k_{L}$ (not shown for brevity).